# Multi: Multimodal Understanding Leaderboard with Text and Images

### v3

@@ 4.2 Settings
We use specialized prompts for each question, an example shown in Figure 5. The prompts are designed carefully according to the features of each type of question and the answer patterns expected. We also modify the input format to fit into official inference guidelines. The complete collection of prompts is presented in C.
@@
For pure-text questions, we use the text as input. For some models like VisCPM, InternVL-1.1, and Yi-VL-34B which compulsorily demand an image in each turn, we feed the model a blank image with color set to RGB(0,0,0) along with plain text in evaluation.
@@ 4.3 Metrics
We focus on objective questions with a certain answer, including multiple-choice and blank-filling questions. We also give a score to each subjective open-ended question based on the similarity to the reference answer. The metrics we use for each type of question:
@@
We follow the metric used in Gaokao and define the total points of an MA question as the number of correct choices, and each correct choice selected is rewarded one point. If the given answer contains any wrong choice, the score will be counted to zero. We report the score ratio (# points / # total points) as the metric. We also report accuracy as a more rigorous metric, where only correctly giving all the choices without wrong ones will be granted points. For example, a question with a correct answer ACE worth 3 points, and the incomplete answer AC will be granted 2 points and the answer BC or ABCE will be granted 0 points. However, on calculating accuracy none will be counted, except for ACE which will be calculated as correct.
@@
We define the total points of a blank-filling question as the number of the blanks marked as [MASK]. It is required in prompts that each line of the given answer corresponds to a blank in order. We follow the most strict standard of exact match. Therefore, only answers exactly matching one of the standard answers will be granted points.
@@
The points and counting method is similar to FB, but we use a loose standard and report normalized ROUGE-L Lin (2004) score for each point. Please note that since the reference answer may be concise or in detail, and there could be other possible answers, the lower score ratio of OP questions is expected and reasonable.
We report the score ratio of all questions as the final metric.
@@ 4.4 Main Experiment Results
In this sub-section we report the performance on Multi as detailed in Table 4, 5, and 6 along with analysis below:
@@
Table 4: The main performance of all models evaluated on Multi. JuH: Level of Junior High School, SeH: Level of Senior High School, Uni: Level of University, Driv: Chinese Driving Test, AAT: Administrative Aptitude Test. We categorized the models into two groups based on whether they are accessed through model weights or via an API service. Models in each group are sorted by their release date. The highest scores in each group are bolded, while the second highest is underlined. The scores higher than the human expert baseline are highlighted.
@@
For questions at the society level, i.e. Driv. and AAT, we anticipate higher scores on the Driving Test. This may be caused by a larger percentage of judgmental questions (in the format of SA with two options), as well as its nature with knowledge of regulations. We are happy to say that Qwen2-VL-72B has surpassed the human expert baseline in this subject.
Furthermore, questions from the Administrative Aptitude Test (AAT), which typically include at least one image and often examine skills in image pattern recognition (illustrated in the first two examples in the left column of Figure 14), tend to have scores around or below randomly choosing baseline. Even the strongest competitor, Qwen2-VL-72B, shows limited success, with a performance of only 34.0% on these questions. This underscores the significant challenge posed by multimodal questions.
Analysis from the perspective of different numbers of image:
In Table 5, we present the performance categorized by image number. While the highest score on the None-image set achieves 86.9% by Qwen2-VL-72B, a significant drop in performance is observed when answering questions with more than one image. Only Gemini 1.5 Pro manages to get a 50.0% score while other strong models are around 40-41%. It is evident that questions requiring more images are more challenging. The high performance on the NI set may also caused by the large portion of easy-driving questions as mentioned before.
@@
In Table 6, we present the performance categorized by question type. A majority of the models achieve their highest scores on the Single Answer Choosing (SA) set, with lower performance on the Multiple Answers Choosing (MA) set. A notable discrepancy is observed between the scores for the MA set and its accuracy, highlighting the smaller models’ inability to identify all correct options accurately. While Qwen2-VL-72B once again beat human expert baseline.
@@
Table 5: Performance on each image type of Multi.
@@
Table 6: Performance on each type of questions of Multi. MA Acc.: Accuracy of MA questions.
@@ 4.5 Is Multi-Elite more challenging?
We conduct evaluations on Multi-Elite, as outlined in Table 7, which includes 500 specifically chosen questions. These questions are selected based on pre-annotated quality and difficulty scores, in addition to the evaluation results on Multi discussed in § 4.4. The selection aims to ensure a more balanced distribution but also picks those high-quality questions that bring challenge to strong MLLMs.
@@
It is important to highlight the accuracy of multiple answers choosing (MA Acc.) as the most demanding task for MLLMs, necessitating a thorough grasp of the relationships between the choices and the questions, and reflecting model reliability of selecting all answers correctly. Meanwhile, the FB set has the largest gap between the highest model performance and human expert baseline, indicating a great challenge for models to answer accurately in these questions.
@@
Table 7: The performance of all models evaluated on Multi-Elite. $\dagger$: these models are chosen to guide the Multi-Elite selection process, thus the performance of these models are relatively lower.
@@ 4.6 Can model utilize Multi-Extend?
The study is conducted on several selected MLLMs, with the prompts for incorporating these knowledge pieces shown in Figure 10. In Figure 6, we present the performance improvement or decline on Multi-Elite with Multi-Extend provided as external knowledge. We also plot the context window size of each model in the figure.
Multi-Extend poses a significant challenge in terms of the necessary window size and the capacity to handle lengthy contexts. It is observed that models equipped with larger window sizes, i.e. GPT-4 series, are more likely to benifit from Multi-Extend, whereas there is a notable decline in performance for MLLMs with smaller window sizes. The increase in tokens may also present a hurdle for models, as those concise questions may become overshadowed or distracted by the extensive context.
4.7 Do these images provides information?
@@
MOSS 36.1 27.3 27.3 ($+$0.0) 27.6 ($+$0.3) - 17.1 20.7 ($+$3.6) 19.0 ($+$1.9) -
DFM-2.0 63.0 28.7 30.2 ($+$1.5) 33.4 ($+$4.7) - 11.3 15.6 ($+$4.3) 14.9 ($+$3.6) -
Qwen2 61.6 35.3 31.6 (--3.7) 35.0 (--0.3) - 18.0 16.2 (--1.8) 17.1 (--0.9) -
InternLM2.5 70.3 39.2 38.5 (--0.7) 38.5 (--0.7) - 24.5 28.6 ($+$4.1) 23.3 (--1.2) -
[…]
VisualGLM 35.1 20.8 21.4 ($+$0.6) 20.4 (--0.4) 25.2 ($+$4.4) 15.3 15.1 (--0.2) 14.5 (--0.8) 9.7 (--5.6)
VisCPM 36.8 27.1 27.6 ($+$0.5) 27.2 ($+$0.1) 28.4 ($+$1.3) 24.8 21.6 (--3.2) 20.9 (--3.9) 16.6 (--8.2)
Chinese-LLaVA 32.3 26.1 26.3 ($+$0.2) 25.5 (--0.6) 22.6 (--3.5) 17.6 19.9 ($+$2.3) 19.6 ($+$2.0) 17.8 ($+$0.2)
Qwen-VL 43.2 30.7 30.3 (--0.4) 31.0 ($+$0.3) 32.7 ($+$2.0) 25.5 25.0 (--0.5) 26.2 ($+$0.7) 20.7 (--4.8)
Yi-VL-34B 63.8 39.9 38.7 (--1.2) 39.4 (--0.5) 42.0 ($+$2.1) 24.1 26.5 ($+$2.4) 24.2 ($+$0.1) 24.5 ($+$0.4)
InternVL-1.1 50.9 33.4 33.3 (--0.1) 33.1 (--0.3) 35.5 ($+$2.1) 24.8 21.9 (--2.9) 22.9 (--1.9) 25.1 ($+$0.3)
InternVL2-8B 78.7 48.4 49.0 ($+$0.6) 48.2 (--0.2) 51.6 ($+$3.2) 32.7 32.6 (--0.1) 28.6 (--4.1) 28.7 (--4.0)
MiniCPM-V-2.6 63.8 37.9 35.0 (--2.9) 38.3 ($+$0.4) 41.8 ($+$3.9) 21.8 25.1 ($+$3.3) 24.0 ($+$2.2) 24.0 ($+$2.2)
Qwen2-VL-7B 78.1 47.8 48.9 ($+$1.1) 48.5 ($+$0.7) 53.7 ($+$5.9) 29.2 30.4 ($+$1.2) 30.5 ($+$1.3) 34.2 ($+$5.0)
Qwen2-VL-72B 86.9 52.1 56.6 ($+$4.5) 56.6 ($+$4.5) 61.5 ($+$9.4) 28.4 33.0 ($+$4.6) 32.0 ($+$3.6) 41.1 ($+$12.7)
[…]
Gemini/Vision Pro 62.5 36.2 36.9 ($+$0.7) 38.4 ($+$2.2) 40.0 ($+$3.8) 18.3 23.2 ($+$4.9) 18.6 ($+$0.3) 24.5 ($+$6.2)
ChatGPT 54.0 6.8 9.9 ($+$3.1) 6.6 (--0.2) - 5.1 10.7 ($+$5.6) 5.5 ($+$0.4) -
GPT-4/V 74.5 11.3 9.7 (--1.6) 1.9 (--9.4) 46.9 ($+$35.6) 8.8 9.4 ($+$0.6) 3.1 (--5.7) 28.1 ($+$19.3)
GPT-4o-mini 67.7 17.5 30.0 ($+$12.5) 29.6 ($+$12.1) 43.0 ($+$25.5) 8.2 14.9 ($+$6.7) 10.3 ($+$2.1) 30.4 ($+$22.2)
Gemini 1.5 Flash 73.8 26.3 38.7 ($+$12.4) 38.7 ($+$12.4) 50.9 ($+$24.6) 11.9 21.4 ($+$9.5) 15.5 ($+$3.6) 40.0 ($+$28.1)
GPT-4o (0806) 80.0 18.5 33.6 ($+$15.1) 33.8 ($+$15.3) 54.1 ($+$35.6) 12.0 17.8 ($+$5.8) 13.2 ($+$1.2) 40.9 ($+$28.9)
Gemini 1.5 Pro 80.2 29.4 43.5 ($+$14.1) 47.6 ($+$18.2) 58.5 ($+$29.1) 11.6 27.9 ($+$16.3) 25.4 ($+$13.8) 50.0 ($+$38.4)
[…]
Table 8: Performance with different image information on Multi. w/o.: No image information is provided. w. cap: Image captions are provided. w. ocr: Image OCR transcripts are provided. w. img: Original images from Multi are provided. We calculate the average improvements for each group of models and different image types.
To assess the necessity of images in Multi for solving problems, we conduct an ablation study where we either remove images from the SI and MI sets or substitute them with textual descriptions, such as captions and OCR-derived text. We utilize BLIP2 Li et al. (2023a) for generating image captions and EasyOCR [^4] https://pypi.org/project/easyocr/ to extract text from images. The results are shown in Table 8.
@@
For questions that involve multiple images (as discussed in the MI column), we categorize models into three groups: 1) Close-source models, specifically GPT-4V and Gemini Pro, which leverage all images and achieve significant improvement. 2) Open-source models capable of handling multiple high-resolution images at a time, namely Qwen2-VL and MiniCPM-V-2.6, which show promising improvements. 3) Open-source models without multi-image support, like VisualGLM, and VisCPM, which show huge performance decline. Only one of the images given could be distracting, especially in a situation where each choice contains an image.
In settings where images are omitted and replaced by their textual descriptions (captions or OCR text), a small improvement is observed with either captions or OCR text for most models. Captions, which generally summarize the images, introduce bilingual elements to the models and usually miss details. OCR text, while detailed, lacks spatial information and is not universally applicable, as some images contain no text at all. Both forms of textual information lower the models’ refusal rate, and LLMs benefit more from these than MLLMs. However, they potentially complicate reasoning processes. Nevertheless, a generic caption is found to more frequently outperform scattered OCR fragments.
@@ 4.8 Will these models refuse to answer if the image is not provided?
(a) SI set without image information
(b) MI set without image information
Figure 7: The portion of answer rejection when evaluated on SI set and MI set without image information as input.
@@ 4.9 Is Multi sensitive to different prompts?
To investigate the sensitivity of Multi to variations in prompt settings, we conducted an ablation study using multiple prompt configurations. Specifically, we examined the following scenarios: (1) base: the original prompt, (2) eng: using English as the prompt language, (3) no bg: removing the question background in the beginning, i.e. You need to use your knowledge of {knowledge} to answer this {question_type} question, (4) cot: adding zero-shot chain-of-thoughtKojima et al. (2022) reasoning instructions. The results are presented in Table 9.
The results indicate that Multi is stable by switching to another language or drop background information, yet will benefit by chain-of-thoughts reasoning.
@@
InternVL2-8B 36.3 35.6 (--0.7) 35.6 (--0.7) 38.3 ($+$2.0)
MiniCPM-V-2.6 21.3 25.0 ($+$3.7) 21.6 ($+$0.3) 26.3 ($+$5.0)
Qwen2-VL-7B 40.0 40.8 ($+$0.8) 39.0 (--1.0) 31.1 (--8.9)
Qwen2-VL-72B 53.1 53.9 ($+$0.8) 54.6 ($+$1.5) 51.1 (--2.0)
GPT-4o-mini 18.5 20.2 ($+$1.7) 21.2 ($+$2.7) 30.7 ($+$12.2)
GPT-4o (1120) 36.3 37.6 ($+$1.3) 38.8 ($+$2.5) 42.0 ($+$5.7)
Gemini 1.5 Flash 33.4 34.0 ($+$0.6) 33.1 (--0.3) 36.4 ($+$3.0)
Gemini 1.5 Pro 46.9 46.1 (--0.8) 45.0 (--1.9) 52.6 ($+$5.7)
[…]
Table 9: The comparison of different prompt settings. base: the original prompt. cot: zero-shot chain-of-thought. eng: English translation of the original prompt. no bg: remove background information.

### v4

@@ 4.2 Settings
We use specialized prompts for each question, an example shown in Figure 5. The prompts are designed carefully according to the features of each type of question and the answer patterns expected. We also modify the input format to fit into official inference guidelines. The complete collection of prompts is presented in D.
@@
For pure-text questions, we use the text as input. For some models like VisCPM, InternVL-1.1, and Yi-VL-34B, which compulsorily demand an image in each turn, we feed the model a blank image with color set to RGB(0,0,0) along with plain text in evaluation.
@@ 4.3 Metrics
We focus on objective questions with a certain answer, including multiple-choice and fill-in-the-blank questions. We also give a score to each subjective open-ended question based on the similarity to the reference answer. We report the score ratio of all questions as the final metric. The metrics we use for each type of question:
@@
We follow the metric used in Gaokao and define the total points of an MA question as the number of correct choices, and each correct choice selected is rewarded one point. If the given answer contains any wrong choices, the score will be counted as zero. We report the score ratio (# points / # total points) as the metric. We also report accuracy as a more rigorous metric, where only correctly giving all the choices without wrong ones will be granted points. For example, a question with a correct answer ACE worth 3 points, and the incomplete answer AC will be granted 2 points, and the answer BC or ABCE will be granted 0 points. However, on calculating accuracy, none will be counted, except for ACE, which will be calculated as correct.
@@
We define the total points of a blank-filling question as the number of the blanks marked as [MASK]. It is required in the prompts that each line of the given answer corresponds to a blank in order. We follow the strictest standard of exact match. Therefore, only answers exactly matching one of the standard answers will be granted points.
@@
The points and counting method is similar to FB, but we use a loose standard and report the normalized ROUGE-L [64] score for each point. Please note that since the reference answer may be concise or in detail, and there could be other possible answers, the lower score ratio of OP questions is expected and reasonable.
@@ 4.4 Main Experiment Results
In this sub-section we report the performance on Multi as detailed in Table 4, 6, and 6 along with analysis below:
@@
Table 4: The main performance (%) of all models evaluated on M ULTI . JuH: Level of Junior High School, SeH: Level of Senior High School, Uni: Level of University, Driv: Chinese Driving Test, AAT: Administrative Aptitude Test. We categorized the models into two groups based on whether they are accessed through model weights or via an API service. Models in each group are sorted by their release date. The highest scores in each group are in bold, while the second highest scores are underlined. The scores higher than the human expert baseline are highlighted.
@@
For questions at the society level, i.e., Driv. and AAT, we anticipate higher scores on the Driving Test. This may be caused by a larger percentage of judgmental questions (in the format of SA with two options), as well as its nature with knowledge of regulations. We are happy to say that Qwen2-VL-72B has surpassed the human expert baseline in this subject.
Furthermore, questions from the Administrative Aptitude Test (AAT), which typically include at least one image and often examine skills in image pattern recognition (illustrated in the first two examples in the left column of Figure 14), tend to have scores around or below the random baseline. Even the strongest competitor, Qwen2-VL-72B, shows limited success, with a performance of only 34.0% on these questions. This underscores the significant challenge posed by multimodal questions.
Analysis from the perspective of different numbers of images:
In Table 6, we present the performance categorized by image number. While the highest score on the None-image set achieves 86.9% by Qwen2-VL-72B, a significant drop in performance is observed when answering questions with more than one image. Only Gemini 1.5 Pro manages to get a 50.0% score while other strong models are around 40-41%. It is evident that questions requiring more images are more challenging. The high performance on the NI set may also be caused by the large portion of easy-driving questions as mentioned before.
@@
In Table 6, we present the performance categorized by question type. A majority of the models achieve their highest scores on the Single Answer Choosing (SA) set, with lower performance on the Multiple Answers Choosing (MA) set. A notable discrepancy is observed between the scores for the MA set and its accuracy, highlighting the smaller models’ inability to identify all correct options accurately. While Qwen2-VL-72B once again beat the human expert baseline.
@@
Table 5: Performance (%) on each image type of M ULTI .
@@
Table 6: Performance (%) on each type of questions of M ULTI . MA Acc.: Accuracy of MA questions.
We keep evaluating new models and updating our leaderboard on our official website  [^4] https://opendfm.github.io/MULTI-Benchmark/#leaderboard. Readers can find other useful links in B.
@@ 4.5 Is M ULTI -E LITE  more challenging?
We conduct evaluations on Multi-Elite, as outlined in Table 7, which includes 500 specifically chosen questions. These questions are selected based on pre-annotated quality and difficulty scores, in addition to the evaluation results on Multi discussed in § 4.4. The selection aims to ensure a more balanced distribution but also picks those high-quality questions that bring a challenge to strong MLLMs.
@@
It is important to highlight the accuracy of multiple answers choosing (MA Acc.) as the most demanding task for MLLMs, necessitating a thorough grasp of the relationships between the choices and the questions, and reflecting model’s reliability of selecting all answers correctly. Meanwhile, the FB set has the largest gap between the highest model performance and human expert baseline, indicating a great challenge for models to answer accurately in these questions.
@@
Table 7: The performance (%) of all models evaluated on M ULTI -E LITE . $\dagger$: these models are chosen to guide the Multi-Elite selection process, thus the performance of these models is relatively lower.
@@ 4.6 Can model utilize M ULTI -E XTEND ?
The study is conducted on several selected MLLMs, with the prompts for incorporating these knowledge pieces shown in Figure 10. In Table 7, we present the performance improvement or decline on Multi-Elite with Multi-Extend provided as external knowledge. We also plot the context window size of each model in the figure.
Multi-Extend poses a significant challenge in terms of the necessary window size and the capacity to handle lengthy contexts. It is observed that models equipped with larger window sizes, i.e., GPT-4 series, are more likely to benefit from Multi-Extend, whereas there is a notable decline in performance for MLLMs with smaller window sizes. The increase in tokens may also present a hurdle for models, as those concise questions may become overshadowed or distracted by the extensive context.
4.7 Do these images provide information?
@@
MOSS 36.1 27.3 27.3 (++0.0) 27.6 (++0.3) - 17.1 20.7 (++3.6) 19.0 (++1.9) -
DFM-2.0 63.0 28.7 30.2 (++1.5) 33.4 (++4.7) - 11.3 15.6 (++4.3) 14.9 (++3.6) -
Qwen2 61.6 35.3 31.6 (-3.7) 35.0 (-0.3) - 18.0 16.2 (-1.8) 17.1 (-0.9) -
InternLM2.5 70.3 39.2 38.5 (-0.7) 38.5 (-0.7) - 24.5 28.6 (++4.1) 23.3 (-1.2) -
Average (-0.73) (++1.00) - (++2.55) (++0.85) -
VisualGLM 35.1 20.8 21.4 (++0.6) 20.4 (-0.4) 25.2 (++4.4) 15.3 15.1 (-0.2) 14.5 (-0.8) 9.7 (-5.6)
VisCPM 36.8 27.1 27.6 (++0.5) 27.2 (++0.1) 28.4 (++1.3) 24.8 21.6 (-3.2) 20.9 (-3.9) 16.6 (-8.2)
Chinese-LLaVA 32.3 26.1 26.3 (++0.2) 25.5 (-0.6) 22.6 (-3.5) 17.6 19.9 (++2.3) 19.6 (++2.0) 17.8 (++0.2)
Qwen-VL 43.2 30.7 30.3 (-0.4) 31.0 (++0.3) 32.7 (++2.0) 25.5 25.0 (-0.5) 26.2 (++0.7) 20.7 (-4.8)
Yi-VL-34B 63.8 39.9 38.7 (-1.2) 39.4 (-0.5) 42.0 (++2.1) 24.1 26.5 (++2.4) 24.2 (++0.1) 24.5 (++0.4)
InternVL-1.1 50.9 33.4 33.3 (-0.1) 33.1 (-0.3) 35.5 (++2.1) 24.8 21.9 (-2.9) 22.9 (-1.9) 25.1 (++0.3)
InternVL2-8B 78.7 48.4 49.0 (++0.6) 48.2 (-0.2) 51.6 (++3.2) 32.7 32.6 (-0.1) 28.6 (-4.1) 28.7 (-4.0)
MiniCPM-V-2.6 63.8 37.9 35.0 (-2.9) 38.3 (++0.4) 41.8 (++3.9) 21.8 25.1 (++3.3) 24.0 (++2.2) 24.0 (++2.2)
Qwen2-VL-7B 78.1 47.8 48.9 (++1.1) 48.5 (++0.7) 53.7 (++5.9) 29.2 30.4 (++1.2) 30.5 (++1.3) 34.2 (++5.0)
Qwen2-VL-72B 86.9 52.1 56.6 (++4.5) 56.6 (++4.5) 61.5 (++9.4) 28.4 33.0 (++4.6) 32.0 (++3.6) 41.1 (++12.7)
Average (++0.29) (++0.40) (++3.08) (++0.69) (-0.08) (-0.18)
Gemini/Vision Pro 62.5 36.2 36.9 (++0.7) 38.4 (++2.2) 40.0 (++3.8) 18.3 23.2 (++4.9) 18.6 (++0.3) 24.5 (++6.2)
ChatGPT 54.0 6.8 9.9 (++3.1) 6.6 (-0.2) - 5.1 10.7 (++5.6) 5.5 (++0.4) -
GPT-4/V 74.5 11.3 9.7 (-1.6) 1.9 (-9.4) 46.9 (++35.6) 8.8 9.4 (++0.6) 3.1 (-5.7) 28.1 (++19.3)
GPT-4o-mini 67.7 17.5 30.0 (++12.5) 29.6 (++12.1) 43.0 (++25.5) 8.2 14.9 (++6.7) 10.3 (++2.1) 30.4 (++22.2)
Gemini 1.5 Flash 73.8 26.3 38.7 (++12.4) 38.7 (++12.4) 50.9 (++24.6) 11.9 21.4 (++9.5) 15.5 (++3.6) 40.0 (++28.1)
GPT-4o (0806) 80.0 18.5 33.6 (++15.1) 33.8 (++15.3) 54.1 (++35.6) 12.0 17.8 (++5.8) 13.2 (++1.2) 40.9 (++28.9)
Gemini 1.5 Pro 80.2 29.4 43.5 (++14.1) 47.6 (++18.2) 58.5 (++29.1) 11.6 27.9 (++16.3) 25.4 (++13.8) 50.0 (++38.4)
Average (++5.96) (++4.00) (++25.13) (++4.72) (-0.34) (++19.15)
Table 8: Performance (%) with different image information on M ULTI . w/o.: No image information is provided. w. cap: Image captions are provided. w. ocr: Image OCR transcripts are provided. w. img: Original images from Multi are provided. We calculate the average improvements for each group of models and different image types.
To assess the necessity of images in Multi for solving problems, we conduct an ablation study where we either remove images from the SI and MI sets or substitute them with textual descriptions, such as captions and OCR-derived text. We utilize BLIP2 [37] for generating image captions and EasyOCR [^5] https://pypi.org/project/easyocr/ to extract text from images. The results are shown in Table 8.
@@
For questions that involve multiple images (as discussed in the MI column), we categorize models into three groups: 1) Closed-source models, specifically GPT-4V and Gemini Pro, which leverage all images and achieve significant improvement. 2) Open-source models capable of handling multiple high-resolution images at a time, namely Qwen2-VL and MiniCPM-V-2.6, which show promising improvements. 3) Open-source models without multi-image support, like VisualGLM, and VisCPM, which show a huge performance decline. Only one of the images given could be distracting, especially in a situation where each choice contains an image.
In settings where images are omitted and replaced by their textual descriptions (captions or OCR text), a small improvement is observed with either captions or OCR text for most models. Captions, which generally summarize the images, introduce bilingual elements to the models, and usually miss details. OCR text, while detailed, lacks spatial information and is not universally applicable, as some images contain no text at all. Both forms of textual information lower the models’ refusal rate, and LLMs benefit more from these than MLLMs. However, they potentially complicate reasoning processes. Nevertheless, a generic caption is found to more frequently outperform scattered OCR fragments.
@@ 4.8 Will these models refuse to answer if the image is not provided?
Figure 7: The portion of answer rejection when evaluated on SI set (left) and MI set (right) without image information as input.
@@ 4.9 Is M ULTI  sensitive to different prompts?
To investigate the sensitivity of Multi to variations in prompt settings, we conducted an ablation study using multiple prompt configurations. Specifically, we examined the following scenarios: (1) BASE: the original prompt, (2) ENG: using English as the prompt language, (3) NO BG: removing the question background in the beginning, i.e., You need to use your knowledge of {knowledge} to answer this {question_type} question, (4) COT: adding zero-shot chain-of-thought[65] reasoning instructions. The results are presented in Table 9.
The results indicate that Multi is stable by switching to another language or dropping background information, yet will benefit from chain-of-thoughts reasoning.
@@
InternVL2-8B 36.3 35.6 (-0.7) 35.6 (-0.7) 38.3 (++2.0)
MiniCPM-V-2.6 21.3 25.0 (++3.7) 21.6 (++0.3) 26.3 (++5.0)
Qwen2-VL-7B 40.0 40.8 (++0.8) 39.0 (-1.0) 31.1 (-8.9)
Qwen2-VL-72B 53.1 53.9 (++0.8) 54.6 (++1.5) 51.1 (-2.0)
GPT-4o-mini 18.5 20.2 (++1.7) 21.2 (++2.7) 30.7 (++12.2)
GPT-4o (1120) 36.3 37.6 (++1.3) 38.8 (++2.5) 42.0 (++5.7)
Gemini 1.5 Flash 33.4 34.0 (++0.6) 33.1 (-0.3) 36.4 (++3.0)
Gemini 1.5 Pro 46.9 46.1 (-0.8) 45.0 (-1.9) 52.6 (++5.7)
Average (++1.0) (++0.4) (++2.9)
Table 9: The comparison of different prompt settings (%). BASE: the original prompt. ENG: English translation of the original prompt. NO BG: remove background information. COT: zero-shot chain-of-thought.